40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Aug 14 04:12:37 2011, Jenny, Update, PSL, Temperature steps and slow actuator railing Picture_6.pngPicture_7.png
    Reply  Sun Aug 14 15:37:39 2011, Jenny, Update, PSL, Temperature steps and slow actuator railing 
Message ID: 5230     Entry time: Sun Aug 14 15:37:39 2011     In reply to: 5228
Author: Jenny 
Type: Update 
Category: PSL 
Subject: Temperature steps and slow actuator railing 

Quote:

Below are some plots from dataviewer of temperature-step data taken over the past 32 hours. (They show minute trends). I am looking at the thermal coupling from the can surrounding the reference cavity on the PSL table to the cavity itself, and trying to measure the cavity temperature response via the control signal sent to heat the NPRO laser, which is locked to the cavity.

Picture_6.png

Picture_7.png

  • Top left: out-of-loop temperature sensor on can surrounding ref cav (RCTEMP)
  • Top right: control signal sent to slow drive of laser (laser heater), which is supposed to follow the cavity temperature (TMP_OUTPUT)
  • Bottom left: in-loop can temperature sensors (MINCOMEAS)
  • Bottom right: room temperature reading (RMTEMP)

 

I stepped the temperature set point from 35 to 36 deg. C for the can at 12:30am last night. Then I waited to see the cavity temperature change and the slow actuator (laser heater: TMP_OUTPUT) follow that change.

I was a bit worried about the oscillations that were occuring in the TMP_OUTPUT signal even long after this temperature step was made, but I figured that they were simply room-temperature changes propagating into the cavity, since they seemed to have a similar pattern to the room-temperature variations, and since it is clear that the out-of-loop temperature sensor on the can (RCTEMP) experiences variations, even when the in-loop sensors are recording no variation.

At 8:46pm tonight I stepped the temperature down 2 degrees to 34 deg. C. The step had a clear effect on TMP_OUTPUT. The voltage to the heater dropped and eventually railed at its lowest output. I'm worried that the loop is unstable, although I haven't ruled out other possibilities, such as that a 2 deg. C temperature step is too large for the loop. I will investigate further in the morning.

 The lock was lost when I came in around noon today to check on it. The slow actuator was still railing.

1) I got lock back for a few minutes, by varying the laser temperature set point manually. TMP_OUTPUT was still reading -30000 cts (minimum allowed) and the transmission was not as high as it had been.

2) I toggled the second filter button off. The TMP_OUTPUT started rising up to ~2000 cts. I then toggled the second filter back on, and TMP_OUTPUT jumped the positive maximum number of counts allowed.

3) I lost the lock again. I turned off the digital output to the slow actuator.

4) I have so far failed at getting the lock back. My main problem is that when the BNC cable to the slow port is plugged in, even when I'm not sending anything to that port, it makes it so that changing the temperature set point manually has almost no effect on the transmission (it looks as though changing the setpoint is not actually changing the temperature, because the monitor shows the same higher order mode even when with +-degree temperature setpoint changes).

ELOG V3.1.3-